A pirate captain in her ship spies her first mate in a dinghy five kilometers away. The pirate captain sails
her ship toward the dinghy at a rate of eight kilometers per hour. The first mate rows his dinghy toward
the pirate ship at a rate of two kilometers per hour. When the captain initially spies the first mate at a
distance of five kilometers, her parrot, Polly, begins flying back and forth between the two at a rate of 40
kilometers per hour. How far does Polly fly in total if she continues her back-and-forth journey until the
pirate ship meets the dinghy?
Answer #1

Distance between ship and dinghy= 5 km.

Speed of ship= 8km/hr.

Speed of dinghy=2 km/hr.

Let the ship and dinghy meet after time t.

In this time, distance travelled by ship=speed *time = 8t.

Similarly,distance travelled by dinghy=speed*time=2t.

Total distance travelled=8t+2t=10t.

So,10t=5 => t= 0.5 hours.

Speed of polly=40 km/hr.

So,distance travelled by polly= speed*time= 40*0.5 = 20 km.